A solid metal cube of side 10cm is heated from 10degree to 60degree,if the linear expansivity of the metal is 1.2 multiplie by 10 raised to power -5 Kelvin raised to power -1, calculate the increase in it's volume.
1
Expert's answer
2018-05-15T09:33:08-0400
V=V_0=(1+α∆T)
∆V=V-V_0=V_0 α∆T=(10)^3×1.2×〖10〗^(-5)×(60-10)=0.6 cm^3